# Frances Dean    Jenna Fields    Radhika Bhalerao    Marie Charpignon    Ahmed Alaa

### v1

@@ Abstract
Instrumental variable (IV) methods mitigate bias from unobserved confounding in observational causal inference but rely on the availability of a valid instrument, which can often be difficult or infeasible to identify in practice. In this paper, we propose a representation learning approach that constructs instrumental representations from observed covariates, which enable IV-based estimation even in the absence of an explicit instrument. Our model (ZNet) achieves this through an architecture that mirrors the structural causal model of IVs; it decomposes the ambient feature space into confounding and instrumental components, and is trained by enforcing empirical moment conditions corresponding to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). Importantly, ZNet is compatible with a wide range of downstream two-stage IV estimators of causal effects. Our experiments demonstrate that ZNet can (i) recover ground-truth instruments when they already exist in the ambient feature space and (ii) construct latent instruments in the embedding space when no explicit IVs are available. This suggests that ZNet can be used as a “plug-and-play” module for causal inference in general observational settings, regardless of whether the (untestable) assumption of unconfoundedness is satisfied.
@@ 1 Introduction
In this paper, we develop a representation learning approach that constructs IVs from observed data by decomposing the input feature space into confounding and instrumental components (Fig. 1). Our model, which we call ZNet, is an encoder architecture that mirrors the structural causal model (SCM) of IVs and is regularized by enforcing the empirical moment conditions that correspond to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). The learned instrumental representation can then be used as input to standard two-stage IV estimators to recover causal effects. When a valid IV is present in the ambient feature space, ZNet recovers a representation that is strongly correlated with it; when no explicit instrument exists, ZNet learns a latent representation that can serve as a pragmatic instrument. The ZNet architecture is compatible with a broad class of downstream IV estimators, and through semi-synthetic experiments we show that learning instrumental representations substantially reduces bias from unobserved confounding. Additionally, we illustrate the learned feature decomposition in real-world, high-dimensional settings using electrocardiogram data.
@@ 2 Preliminaries
Thus, with the instrument $Z$, we can estimate the counterfactual $\mathbb{E}[Y|do(T),X]$ by estimating the two observable functions $\mathbb{E}[Y|X,Z]$ and $F(T|X,Z)$. While this inverse problem is ill-posed, it provides a practical framework for estimating counterfactuals, and identification is possible under certain conditions (Newey and Powell, 2003). [^1] For example, in the linear case, Two-Stage Least Squares Regression (TSLS) allows for identification of causal effects. A two-stage regression fits a model $\widehat{F}(T|X,Z)$, and estimates $\mathbb{E}[Y|do(T),X]$ by replacing $F(T|X,Z)$ with $\widehat{F}(T|X,Z)$ in (2). With an estiimate of $\mathbb{E}[Y|do(T),X]$, we can derive the conditional average treatment effects (CATE) and average treatment effects (ATE):
@@ 3 Learning Instruments from Data
Moment conditions on $f(.)$ and $g(.)$. A valid decomposition of the observed features $X$ into confounding and instrumental components must induce a factorization of the joint distribution $\mathbb{P}(g(X),f(X),U,T,Y)$ that is consistent with the SCM defined by (1) and (2). In particular, the relevance condition requires that the instrumental component $g(X)$ be predictive of the treatment, i.e., $\operatorname{Cov}(g(X),T)\neq 0$. Exclusion restriction requires that all direct effects of $X$ on the outcome $Y$ be mediated through the confounding component $\widetilde{X}=f(X)$; equivalently, $\operatorname{Cov}(f(X),Y)\neq 0$ while $\operatorname{Cov}(f(X),g(X))=0$. As we discuss later in detail in Section 5, we enforce these covariance relationships through the architecture and loss function of our proposed model.
The covariance structure described above ensures that the learned instrument $\widetilde{Z}$ will be unconfounded and independent of the error $e_{Y}$ as long as $X$ is unconfounded by $U$. The assumption that $X$ is not influenced by $U$ is standard to allow for classical IV regression and straightforward IV generation (Yuan et al., 2022; Li and Yao, 2024; Cheng et al., 2023; Chou et al., 2024). Since our feature decomposition produces both a confounder and an instrument, we add an additional constraint based on the following Lemma to ensure that $\widetilde{Z}$ is valid even when $X$ may be influenced by $U$.
Lemma 1. If $Z\sim\mathcal{N}(0,\sigma^{2})$, and suppose that the random variables $(Z,e_{Y},X,T)$ satisfy $\mbox{Cov}(Z,e_{Y}-\mathbb{E}[e_{Y}|X,T])=0$. Then it follows that $\mbox{Cov}(Z,e_{Y})=0$.
Proof. Notice that as $\mathbb{E}[Z]=0$, we have
$0$ $=\mbox{Cov}(Z,e_{Y}-\mathbb{E}[e_{Y}|X,T])$
$=\mathbb{E}[Z(e_{Y}-\mathbb{E}[e_{Y}|X,T])]-\mathbb{E}[Z]\,\mathbb{E}[(e_{Y}-\mathbb{E}[e_{Y}|X,T])]$
$=E[Z(e_{Y}-\mathbb{E}[e_{Y}|X,T])]$
$=\mathbb{E}[Z\cdot e_{Y}]-\mathbb{E}[Z]\mathbb{E}[e_{Y}|X,T]=\mbox{Cov}(Z,e_{Y}).\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\square$
This additional covariance term can be evaluated by fitting a regression model for $\hat{Y}=\mathbb{E}[Y|X,T]$ and then computing its residuals as $Y-\mathbb{E}[Y|X,T]$. Notice that $Y-\mathbb{E}[Y|X,T]=Y-\varphi(X,T)-(E[Y|X,T]-\varphi(X,T))=e_{Y}-\mathbb{E}[e_{Y}|X,T]$. Regardless of whether any existing instruments are normally distributed, if we can construct $g(X)=\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2})$ to have no covariance with the residuals $e_{Y}-\mathbb{E}[e_{Y}|X,T]$, i.e. with $\tilde{\varepsilon}_{Y}:=Y-\mathbb{E}[Y|X,T]$, then $\widetilde{Z}$ has zero covariance with $e_{Y}$ by Lemma 1. To sum up, $f$ and $g$ need to satisfy the following moment restrictions:
@@
Moment Condition 3: $\mbox{Cov}(f(X),Y)>0$.
Moment Condition 4: $\mbox{Cov}(T,g(X))>0$.
@@ 5.3 ZNet Implementation and Training Details
$+\alpha_{3}(1-\widehat{\operatorname{Cov}}(\varphi(f(X),T),Y)^{2})$
$+\alpha_{4}(1-\widehat{\operatorname{Cov}}(\pi(g(X)),T)^{2})$
$+\alpha_{5}\widehat{\operatorname{Cov}}(g(X),\tilde{\varepsilon}_{Y})^{2}+\alpha_{6}\widehat{\operatorname{Cov}}(\pi(g(X)),T)^{2}$
@@
Hyperparameter tuning. All hyperparameters, including all loss term weights and whether the moment constraints are approximated via Pearson correlations or MI, were tuned using Bayesian optimization implemented in Botorch (Balandat et al., 2020). We perform optimization corresponding to the two stages of IV regression. For each IV generation, we maximize the instrument’s relevance F-Statistic and minimize the correlation between learned $\tilde{X}$ and using Botorch’s native adaptation of the Noisy Expected Improvement acquisition function for multi-objective optimization. We then choose the parameter set from the Pareto front with the highest F-Statistic. We tune the treatment effect estimators to simultaneously minimize the MSE of the model’s ATE against a nearest-neighbors (NN) ATE and the MSE of estimated factual $Y$, again with the Noisy Expected Improvement acquisition function. The parameter set is selected from the Pareto front by the least MSE of NN.
@@ 6.3 Application of ZNet to Unstructured Data
ZNet in the unstructured data setting is particularly powerful because high dimensional data can contain additional information recoverable as an instrument more readily. We believe this is especially impactful in the setting of unstructured health data. We demonstrate the ability of ZNet to recover instruments in such settings using a semi-synthetic electrocardiogram (ECG) dataset. Data from 35,463 ECGs from a public Physionet dataset serve as $X$, age and sex serve as hidden confounders $U$, and variables $T$ and $Y$ are sampled, analogous to the process for IHDP data, in this case as a random linear combination of eight interpretable, derived features of the ECG, i.e. heart rate, QT interval, QRS amplitude, etc. (Zheng et al., 2022, 2000; Goldberger et al., 2000). As with most unstructured health data, no true instrument exists. Despite that, ZNet still recovers an instrument representation and more accurately predicts treatment effects than ordinary least squares regression using the tabular features (Table 5). Here, we use ZNet with Deep IV since it is consistently more accurate than DF-IV.
@@ 7 Discussion
The data generation process of any real world dataset is untestable. We can never confirm a lack of unobserved confounding, even when evaluating a candidate instrument. ZNet’s strong performance across settings suggests that it can serve as a “plug-in" causal inference estimator for causal inference in general observational settings, regardless of whether the untestable assumption of unconfoundedness is satisfied. Our evaluations show ZNet to be particularly well suited when unobserved confounders influence the observed data and where instrument candidates do not exist. These are the most challenging and general settings.
@@
Solutions to the ZNet loss minimization problem always give a representation that serves empirically as an instrument since empirical IV constraints are explicitly embedded in the loss function. This instrument can then be used in any downstream IV regression where satisfying the standard IV criteria (or, equivalently, ZNet moment constraints) implies the validity of subsequent causal inference. The strength and validity of ZNet generated instruments vary in finite settings. Variance in finite samples affects our ability to predict treatment effects. However, IV estimation is also limited in general by a lack of theoretical guarantees of identifiably without conditions on the SCM. This theoretically limits not only our approach but IV estimation in general. Still, strong empirical results and opportunities in high dimensional data suggest the value in the increased use of these methods beyond linear settings.

### v2

# Causal Effect Estimation with Learned Instrument Representations

@@ Abstract
Instrumental variable (IV) methods mitigate bias from unobserved confounding in observational causal inference but rely on the availability of a valid instrument, which can often be difficult or infeasible to identify in practice. In this paper, we propose a representation learning approach that constructs instrumental representations from observed covariates, which enable IV-based estimation even in the absence of an explicit instrument. Our model (ZNet) achieves this through an architecture that mirrors the structural causal model of IVs; it decomposes the ambient feature space into confounding and instrumental components, and is trained by enforcing empirical moment conditions corresponding to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). Importantly, ZNet is compatible with a wide range of downstream two-stage IV estimators of causal effects. Our experiments demonstrate that ZNet can (i) recover ground-truth instruments when they already exist in the ambient feature space and (ii) construct latent instruments in the embedding space when no explicit IVs are available. Our work suggests when ZNet can be used as a module for causal inference in general observational settings.
@@ 1 Introduction
In this paper, we develop a representation learning approach that constructs IVs from observed data by decomposing the input feature space into confounding and instrumental components (Fig. 1). Our model, which we call ZNet, is an encoder architecture that mirrors the structural causal model (SCM) of IVs and is regularized by enforcing the empirical moment conditions that correspond empirically to the defining properties of valid instruments (i.e., relevance, exclusion restriction, and instrumental unconfoundedness). The learned instrumental representation can then be used as input to standard two-stage IV estimators to recover causal effects. When a valid IV is present in the ambient feature space, ZNet recovers a representation that is strongly correlated with it; when no explicit instrument exists, ZNet learns a latent representation that can serve as a pragmatic instrument. The ZNet architecture is compatible with a broad class of downstream IV estimators, and through semi-synthetic experiments, we show that learning instrumental representations substantially reduces bias from unobserved confounding. Additionally, we illustrate the learned feature decomposition in real-world, high-dimensional settings using electrocardiogram data.
@@ 2 Preliminaries
Thus, with the instrument $Z$, we can estimate the counterfactual $\mathbb{E}[Y|do(T),X]$ by estimating the two observable functions $\mathbb{E}[Y|X,Z]$ and $F(T|X,Z)$. While this inverse problem is ill-posed, it provides a practical framework for estimating counterfactuals, and identification is possible under certain conditions (Newey and Powell, 2003). [^1] For example, in the linear case, Two-Stage Least Squares Regression (TSLS) allows for identification of causal effects. A two-stage regression fits a model $\widehat{F}(T|X,Z)$, and estimates $\mathbb{E}[Y|do(T),X]$ by replacing $F(T|X,Z)$ with $\widehat{F}(T|X,Z)$ in (2). With an estimate of $\mathbb{E}[Y|do(T),X]$, we can derive the conditional average treatment effects (CATE) and average treatment effects (ATE):
@@ 3 Learning Instruments from Data
Moment conditions on $f(.)$ and $g(.)$. A valid decomposition of the observed features $X$ into confounding and instrumental components must induce a factorization of the joint distribution $\mathbb{P}(g(X),f(X),U,T,Y)$ that is consistent with the SCM defined by (1) and (2). We consider constructing these conditions using empirical analogues for their measurement. In particular, the relevance condition suggests that the instrumental component $g(X)$ be predictive of the treatment, i.e., $\operatorname{Cov}(g(X),T)\neq 0$. Exclusion restriction requires that all direct effects of $X$ on the outcome $Y$ be mediated through the confounding component $\widetilde{X}=f(X)$; i.e., $\operatorname{Cov}(f(X),Y)\neq 0$ while $\operatorname{Cov}(f(X),g(X))=0$. We an additional constraint on the model residuals to encourage exogeneity, i.e. that $\operatorname{Cov}(g(X),\tilde{\varepsilon}_{Y})=0$. As we discuss later in detail in Section 5, we enforce these covariance relationships through the architecture and loss function of our proposed model. We remark that these conditions are only equivalent to the IV conditions under the case that all variables are Guassian or when comparing variables using mutual information rather than covariance, thus our work is not theoretical but empirical in nature.
The learned instrument $\widetilde{Z}$ will be uncorrelated with the error $e_{Y}$ as long as $X$ is unconfounded by $U$ since it is derived directly from $X$. The assumption that $X$ is not influenced by $U$ is standard to allow for classical IV regression and straightforward IV generation (Yuan et al., 2022; Li and Yao, 2024; Cheng et al., 2023; Chou et al., 2024). To explore capabilities beyond this strong assumption, we include experiments with data that has confounding of $X$ by $U$. To sum up, we desire $f$ and $g$ to satisfy the following moment restrictions to learn an empirical approximation of an IV:
@@
Moment Condition 3: $\mbox{Cov}(f(X),Y)\neq 0$.
Moment Condition 4: $\mbox{Cov}(T,g(X))\neq 0$.
@@ 5.3 ZNet Implementation and Training Details
$+\alpha_{3}(-\widehat{\operatorname{Cov}}(f(X),Y)^{2})$
$+\alpha_{4}(-\widehat{\operatorname{Cov}}(g(X),T)^{2})$
$+\alpha_{5}\widehat{\operatorname{Cov}}(g(X),\tilde{\varepsilon}_{Y})^{2}+\alpha_{6}\widehat{\operatorname{Cov}}(g(X),f(X))^{2}$
@@
Hyperparameter tuning. All hyperparameters, including all loss term weights and whether the moment constraints are approximated via Pearson correlations or MI, were tuned using Bayesian optimization implemented in Botorch (Balandat et al., 2020). We perform optimization corresponding to the two stages of IV regression. For each IV generation, we maximize the instrument’s relevance F-Statistic and minimize the correlation between learned $\tilde{X}$ and Z{Z} using Botorch’s native adaptation of the Noisy Expected Improvement acquisition function for multi-objective optimization. We then choose the parameter set from the Pareto front with the highest F-Statistic. We tune the treatment effect estimators to simultaneously minimize the MSE of the model’s ATE against a nearest-neighbors (NN) ATE and the MSE of estimated factual $Y$, again with the Noisy Expected Improvement acquisition function. The parameter set is selected from the Pareto front by the least MSE of NN.
@@ 6.3 Application of ZNet to Unstructured Data
ZNet in the unstructured data setting is particularly powerful because high dimensional data can contain additional information recoverable as an instrument more readily. We believe this is especially impactful in the setting of unstructured health data. We demonstrate the ability of ZNet to recover instruments in such settings using a semi-synthetic electrocardiogram (ECG) dataset. Data from 35,463 ECGs from a public Physionet dataset serve as $X$, age and sex serve as hidden confounders $U$, and variables $T$ and $Y$ are sampled, analogous to the process for IHDP data, in this case as a random linear combination of eight interpretable, derived features of the ECG, i.e. heart rate, QT interval, QRS amplitude, etc. Zheng et al. (2022, 2000); Goldberger et al. (2000). As with most unstructured health data, no true instrument exists. Despite that, ZNet still recovers an instrument representation and more accurately predicts treatment effects than ordinary least squares regression using the tabular features (Table 5). Here, we use ZNet with Deep IV since it is consistently more accurate than DF-IV.
@@ 7 Discussion
The data generation process of any real world dataset is untestable. We can never confirm a lack of unobserved confounding, even when evaluating a candidate instrument. ZNet’s strong performance across settings suggests that it can serve to mitigate unobserved confounding in general observational settings, regardless of whether the untestable assumption of unconfoundedness is satisfied. Our evaluations show ZNet to be particularly well suited when unobserved confounders influence the observed data and where instrument candidates do not exist. These are the most challenging and general settings. ZNet does not, however, guarantee empirically in all cases that instruments are valid, and our work does not prove that ZNet representations are IVs.
@@
Solutions to the ZNet loss minimization problem give a representation that serves empirically as an instrument since empirical analogues of IV constraints are embedded in the loss function. This instrument can then be used in any downstream IV regression where satisfying the standard IV criteria implies the validity of subsequent causal inference. The strength and validity of ZNet generated instruments vary in finite settings. Variance in finite samples affects our ability to predict treatment effects. IV estimation is also limited in general by a lack of theoretical guarantees of identifiably without conditions on the SCM. This theoretically limits not only our approach but IV estimation in general. Moreover, we emphasize that our IV representation method does not and cannot mitigate unobserved confounding in all settings. To produce an empirically valid instrument, ZNet requires assuming either that the observed data are not confounded or the conditions of Lemma 1 are satisfied. Still, our experiments and opportunities in high dimensional data suggest the value in the increased research into the use of these methods beyond linear settings.